Adv. Math. Econ. 11, 105-116 (2008)

Rational expectations can preclude trades

Takashi Matsuhisa^{1} and Ryuichiro Ishikawa^{2}

1 Department of Natural Sciences, Ibaraki National College of Technology, 866 Nakane, Hitachinaka-shi, Ibaraki 312-8508, Japan

2 Graduate School of Systems and Information Engineering, University of Tsukuba, 1-1-1, Ten-nodai, Tsukuba, Ibaraki 305-8573, Japan

Received: June 4, 2007 Revised: October 2, 2007

JEL classification: D51, C78, D61

Mathematical Subject Classification (2000): 91A10, 91B44, 91B50

Abstract. We reconsider the no trade theorem in an exchange economy where the traders have non-partition information. By introducing a new concept, rationality of expectations, we show some versions of the theorem different from previous works, such as Geanakoplos (1989). We also reexamine a standard assumption of the no trade theorem: the common prior assumption.

Key words: No trade theorem, Ex ante Pareto optimum, Common knowledge, Rational expectations equilibrium